OpenStudy (anonymous):

Determine if the following is always, sometimes, or never true. y = -1/5 x + 3 is a function. Can someone explain this to me plz

OpenStudy (anonymous):

Informally speaking, a function is a rule that associates one element from one set with a unique element in another set.

OpenStudy (aravindg):

yep this is always a function

OpenStudy (aravindg):

you see for a relation to become a function it needs to just have 2 properties: 1.every element in domain has an image 2. there is only one image for each x value

OpenStudy (aravindg):

clearly y = -1/5 x + 3 gives us a y value for all x inputs and there will be only one image for each x value

OpenStudy (aravindg):

infact its a straight line function with slope -1/5

OpenStudy (aravindg):

hope that helps

OpenStudy (anonymous):

Thanks!

OpenStudy (aravindg):

yw :)

OpenStudy (whpalmer4):

If you are graphing the equation, look at the graph and see if there is any point at which a vertical line would cross the graph twice at the same value of x. If so, it is not a function. So y = x^2 is a parabola opening upward, and no vertical line crosses it more than once for a given value of x. However, x = y^2 is the same parabola except opening to the right, and for any positive value of x, a vertical line crosses the parabola twice, so it is not a function.